The number of people n in the mall is modeled by the formula: n=-100x(sq)+1200x where x is the number of hours since the mall oopened this morning. If the mall opened at 9 am, when were there 2000 or more people in the mall?
Answer by hydromojo(20)


# of people = -100x^2 + 1200x
for 2000 people to be there,
2000 = 100x^2 + 1200x
re-arranging,
-100x^2 + 1200x = 2000
i.e. -x^2 + 12x = 20
-x^2 +12x -20 = 0
-x^2 + 10x + 2x - 20 = 0
solve, you get x = 2hrs